An isosceles triangle has legs that are 15 cm and a perimeter of 48 cm. What is the perimeter of a similar isosceles triangle with a base of 12 cm?
Answer by edjones(8007)

15+15+18=48
12/18=2/3
15 * 2/3 = 10
10+10+12= 32 cm Perimeter.
.
Ed